# Predicting Field Experiments with Large Language Models

### v1

@@ 6 Conclusions
In this paper, we proposed an LLM-powered framework that automatically extracts information from existing papers and predicts field experimental conclusions. To the best of our knowledge, this is the first paper to provide an automated framework for predicting such conclusions. Rather than merely introducing the framework, this is also the first work that examines its fidelity on a large scale of samples and clearly delineates infeasible scenarios in the experimental simulation literature. Building on these foundational efforts, our new results indicate that incorporating a CoT strategy generally enhances predictive performance, whereas merely increasing the number of repetitions does not produce a significant effect. Furthermore, iterative refinements to LLMs consistently improve performance, although the rate of improvement diminishes over time. Notably, the performance gains attributable to CoT appear to be model-sensitive. Taken together, these findings underscore the potential of the LLM-driven frameworks in advancing automated predictions for field experiments while also clarifying the practical constraints that guide their effective use.

### v2

@@ 6. Conclusion
In this paper, we propose an LLM-powered framework that automatically extracts information from existing papers and predicts field experimental conclusions. To the best of our knowledge, this is the first paper to provide an automated framework for predicting such conclusions for field experiments. Rather than merely introducing the framework, our work also examines its fidelity on a large scale of samples and achieves a considerable accuracy of 78%. Furthermore, the paper discovers that incorporating a CoT strategy generally enhances predictive performance in this scenario, whereas the performance gains attributable to CoT appear to be model-dependent. Furthermore, iterative refinements to LLMs consistently improve performance, although the rate of improvement diminishes over time.
Interestingly, the paper also finds that the distributions of prediction accuracy are either bimodal or negatively skewed, with a large number of samples concentrating on the two extremes. To explore this phenomenon, the paper regresses conclusion prediction accuracy on topic components and sentiments, revealing that interaction effects between certain topics and sentiments could affect the LLMs’ prediction performance in this task.
Taken together, these findings underscore the potential of the LLM-driven frameworks in advancing automated predictions for field experiments while also clarifying the practical constraints that guide their effective use.
{strip}
Table 3. OLS Regression Results for Conclusion Accuracy
(1) (2) (3) (4) (5) (6)
Variable GPT-4o Basic GPT-4o CoT GPT-4 Turbo Basic GPT-4 Turbo CoT GPT-3.5 Turbo Basic GPT-3.5 Turbo CoT
Constant 0.919 (0.721) -0.240 (0.929) 3.397 (0.249) 1.860 (0.498) 4.026 (0.182) 2.820 (0.218)
Gender -0.138 (0.957) 1.006 (0.710) -2.932 (0.320) -1.084 (0.693) -3.628 (0.230) -2.245 (0.327)
Ethnicity -0.342 (0.893) 1.103 (0.679) -2.903 (0.318) -1.347 (0.620) -3.485 (0.242) -2.209 (0.328)
Social Norms -0.124 (0.962) 1.129 (0.677) -2.483 (0.402) -1.241 (0.653) -3.607 (0.235) -2.251 (0.328)
Ethical Dilemmas -0.729 (0.779) 0.908 (0.738) -2.821 (0.341) -0.600 (0.828) -3.156 (0.298) -2.068 (0.369)
Age -0.151 (0.954) 0.838 (0.757) -2.868 (0.332) -1.233 (0.655) -3.366 (0.267) -2.231 (0.332)
Socioeconomic Status -0.006 (0.998) 1.071 (0.692) -2.747 (0.352) -1.108 (0.687) -3.486 (0.249) -2.120 (0.355)
Other Topics -0.218 (0.933) 0.955 (0.724) -2.880 (0.329) -1.263 (0.646) -3.423 (0.257) -2.225 (0.331)
Gender × Favorability -0.031 (0.897) 0.478* (0.057) 0.125 (0.647) 0.089 (0.727) -0.313 (0.264) 0.202 (0.341)
Ethnicity × Sentiment 0.349** (0.027) 0.242 (0.142) 0.519*** (0.004) 0.439*** (0.009) 0.126 (0.493) 0.323** (0.021)
Social Norms × Sentiment 0.112 (0.615) 0.121 (0.602) 0.191 (0.453) 0.355 (0.134) 0.692*** (0.008) 0.333* (0.092)
Ethical Dilemmas × Sentiment -0.605** (0.028) -0.318 (0.269) -0.544* (0.083) -0.810*** (0.006) -0.525 (0.103) -0.346 (0.156)
Age × Sentiment 0.225 (0.247) -0.187 (0.357) 0.447** (0.044) 0.176 (0.394) 0.114 (0.617) 0.122 (0.479)
Socioeconomic Status × Sentiment -0.011 (0.893) -0.082 (0.345) 0.101 (0.289) 0.124 (0.161) 0.051 (0.599) 0.062 (0.401)
Other Topics × Sentiment 0.126* (0.053) 0.115* (0.091) 0.081 (0.275) 0.199*** (0.004) -0.019 (0.800) 0.126** (0.029)
$R^{2}$ 0.041 0.017 0.065 0.076 0.033 0.079
Adj. $R^{2}$ 0.027 0.002 0.051 0.062 0.019 0.065
F-statistic 2.859 1.138 4.660 5.542 2.323 5.769
Prob. (F-statistic) 0.0003 0.319 0.000 0.000 0.004 0.000
Observations 955 955 955 955 955 955
Note: Coefficients with p-values in parentheses. *** $p<0.01$, ** $p<0.05$, * $p<0.1$. Also, for interaction terms, since our definition of pro-female content includes negative sentiment when the context is detrimental to the male, it’s not sufficient to represent the gender favorability label with regular sentiment label. Therefore, for the gender topic, we explored its interaction with gender favorability label, while for the rest of the topics, we explored their interactions with sentiment.
Appendix A Data
Figure 7. A Summary of Qualified Papers.
[…]
A bar chart showing the number of selected field experiment papers across different years or journals, highlighting the temporal or disciplinary distribution of the 276 papers used.
Appendix B Prompt Templates
B.1. The Prompt for Extractions
Figure 8. The Prompt for Extractions.
[…]
Screenshot of the prompt used to extract key field experiment settings and conclusions using Claude.
B.2. The Prompt for Variant Generation
Figure 9 shows the extraction prompt used to retrieve experiment settings and conclusions from academic papers using Claude.
Figure 9. The Variant Generation Process.
[…]
Visual showing how reversed and unrelated variants are generated from the original experimental conclusion using LLM prompts.
B.3. Prompts for Prediction
Figure 10. The Basic Prediction Prompt.
[…]
Prompt template used to perform basic prediction among multiple conclusions in a field experiment.
Figure 11. The Chain-of-Thought (CoT) Prediction Prompt.
[…]
Chain-of-thought prediction prompt where the model is instructed to reason through the decision before choosing a conclusion.
B.4. Prompts for Topics and Sentiments Labeling
Figure 12. Topics Labeling Prompt.
[…]
Prompt used to assign topic relevance scores (e.g., gender, age, ethnicity) to each experimental context.
Figure 13. Sentiment Labeling Prompt.
[…]
Prompt used to classify the sentiment (positive, neutral, negative) associated with a conclusion.
Figure 14. Gender Favorability Labeling Prompt.
[…]
Prompt used to label whether a conclusion’s framing is favorable or unfavorable toward a specific gender.
Appendix C Regression Variables
Table 4. Regression Variables
Variable Description
Conclusion Accuracy (DV) Conclusion accuracy defined in Equation 1.
Gender Percentage indicating how strongly the context is associated with gender, obtained via labeling in Figure 12.
Ethnicity Same as above, for ethnicity.
Social Norms Same as above, for social norms.
Ethical Dilemmas Same as above, for ethical dilemmas.
Age Same as above, for age-related context.
Socioeconomic Status Same as above, for socioeconomic status.
Other Topics Same as above, for all remaining topics.
Sentiment Sentiment score: 1 (positive), 0 (neutral), -1 (negative), labeled via Figure 13.
Gender Favorability Gender-specific score: 1 (pro-female or anti-male), 0 (neutral), -1 (pro-male or anti-female), from Figure 14.